OpenStudy (anonymous):

Solve the equation for the given variables 7-6a=6-7a

OpenStudy (anonymous):

7-6a=6-7a/a = -1

OpenStudy (precal):

7-6a=6-7a add 7a to both sides 7+a=6 -7 to both sides a=-1